Determine the profit-maximizing prices when a firm faces two markets where the inverse demand curves are

Market​ A: Pa=80-2Qa

where demand is less​ elastic, and

Market​ B: Pb=60-1Qb

where demand is more​ elastic, and

Marginal Cost = m=40

for both markets.

For Market A Pa= $___

For Market B Pb = $___
Answer #1

Market B г. - 0 - 1 3, Твь с Gosp - sp? MR b = 60 - 2 db Al-equilisunen 6 - 2gb Чо 2 cg, c 2 O C) - to Pь с 60 — 1С со) a soMask - А Pa = 80 - 200 TRa a Pada 2 8oqa - 22 MRa a dCTRa) DOA - Во - ча та - чо At equliberum mRame $o - ча чо Чga - ЧО Sa 2

Pa=60

Pb=50